If I'm traveling at 215 miles and hour how long will it take to go 38 miles?
Found 3 solutions by richard1234, rfer, nyc_function:
Answer by richard1234(5390)

D = rt

38 mi = (215 mi/hr)t

t = 38/215 hr = .1767 hr = 10.60 min

Answer by rfer(12730)

215/1=38/x
215x=38
x=38/215
x=0.177 hrs
---------------
x=60*0.177
x=10.6 min

Answer by nyc_function(2733)






We use the formula D = rate times time.

We write the formula this way: D = rt.

Let D = 38 miles

Let r = 215 miles per hour

We need to find t.

D = rt

38 = 215t

38/215 = t

0.17674418604 = t

We now convert 0.17674418604 to time in hours, minutes, and seconds.

It will take you 0.0 hours, 10.0 minutes, 36.0 seconds.

To convert 36.0 seconds to minutes, divide 36 by 60 seconds in one minute.

This division will yield the decimal answer 0.6.

Adding 0.6 to 10.0 gives us 10.60 minutes.

Remember, you are moving at a speed of 215 miles per hour.

The answer is 10.60 minutes.

I hope this helps.

If you have a fax number, I can fax my answers to you for other questions that you may have. It is easier to write on paper and then fax my answers to students. If you have a fax number, send it to: nycfunction@yahoo.com
Mr. Feliz